OpenStudy (anonymous):

If the ink in the period at the end of this sentence weighs 1.3 x 10^-8 pounds, and if there are 30000 peroids printed in this book, then how much does this ink weigh together? Express each number in scientific Notation and perform the calculation in scienctific notation.

OpenStudy (unklerhaukus):

what kind of computer screen has ink?

OpenStudy (unklerhaukus):

\[1.3 \times 10^{-8} (Lb~of~ink/period) \times 30000(period)\] \[=3.9\times 10^{-4} (Lb~ink)\]

OpenStudy (anonymous):

I dont know what type of computer does, but thanks for your answers!!

OpenStudy (unklerhaukus):

i don't know if pixels weigh more when dark or bright